One person walked to 'B', a distance of 21m from 'A'. He then rounded the circle with 'A' as the focal point and returned again to 'A'. So how many total runs did I do?
1
Expert's answer
2019-08-13T11:19:34-0400

The total distance is the sum of three distances: from A to B, circular motion and from B to A:


"L=L_{AB}+L_{circ}+L_{BA}"

As the distance from A to B is 21 meter, the radius of the circle is also 21 meter.

So:


"L=21+2*\pi*21+21=174 \; m"

Answer: Total distance is 174m